l(17)___________m________n(95)
if m is a point 2/3 of the way from l and n on the number line , the number located at m is
a 11 b 43 c 56 d 69 e 78

Answer by stanbon(75887)

l(17)___________m________n(95)
if m is a point 2/3 of the way from l and n on the number line , the number located at m is:
------
m = 17 + (2/3)(95-17)
m = 17 + (2/3)(78)
m = 17 + 52
m = 69
================
Cheers,
Stan H.
================

a 11 b 43 c 56 d 69 e 78